OpenStudy (anonymous):

Solve the equations and check your solutions. If there is no solution say so 10/3x-8=2/x need the steps please thanks in advance

OpenStudy (anonymous):

is it 10/3x -8 or 10\(3x-8)?

OpenStudy (anonymous):

10/3x=2/x+8 10/3x=(2+8x)/x 3x(2+8x)=10x 6x+24x^2=10x 6x+24x^2-10x=0 24x^2-4x=0 4x(8x-1)=0 x=0 or x=1/8